# Risk Bounds for Mixture Density Estimation on Compact Domains via the $h$-Lifted Kullback–Leibler Divergence

### v1

@@ 6 Numerical experiments
In this section, we discuss the computability and computation of $\textrm{KL}_{h}$ estimation problems and provide empirical evidence towards the rates obtained in Theorem 9. Namely, we seek to develop a methodology for computing $h$-MLLEs, and to use numerical experiments to demonstrate that the sequence of expected $h$-lifted KL divergences between some density $f$ and a sequence of $k$-component mixture densities from a suitable class $\mathcal{P}$, estimated using $n$ observations does indeed decrease at rates proportional to $1/k$ and $1/\sqrt{n}$, as $k$ and $n$ increase.
@@ 6.1 Minorization–Maximization algorithm
One solution for computing problems of kind (4) is to employ an MM algorithm. To do so, we first write the objective of (4) as
@@
$\lim_{s\rightarrow\infty}\inf_{\psi_{k}^{*}\in\Psi_{k}^{*}}\left\|\psi_{k}^{\left(s\right)}-\psi_{k}^{*}\right\|_{2}=0$
where,
$\Psi_{k}^{*}=\left\{\psi_{k}^{*}\in\Psi_{k}:\left.\frac{\partial L_{h,n}}{\partial\psi_{k}}\right|_{\psi_{k}=\psi_{k}^{*}}=0\right\}.$
@@ 6.2 Experimental setup
where $\chi_{\mathcal{A}}$ is the characteristic function that takes value $1$ if $x\in\mathcal{A}$ and 00, otherwise. Note that neither $f_{1}$ nor $f_{2}$ are in $\mathcal{C}$. In particular, $f_{1}\left(x\right)=0$ when $x\in\left(\nicefrac{{2}}{{5}},\nicefrac{{3}}{{5}}\right)$, and $f_{2}(x)=0$ when $x=1/2$, and hence neither densities are bounded away from 0, on $\mathcal{X}$. Thus, the theory of Rakhlin et al. (2005) cannot be applied to provide bounds for the expected KL divergence between MLEs of beta mixtures and the pair of targets. We visualize $f_{1}$ and $f_{2}$ in Figure 1.
@@
To observe the rate of decrease of the $h$-lifted KL divergence between the targets and respective sequences of $h$-MLLEs, we conduct two experiments E1 and E2. In E1, our target density is set to $f_{1}$ and $h_{1}=\beta\left(\cdot;1/2,1/2\right)$. For each $n\in\left\{2^{10},\dots,2^{15}\right\}$ and $k\in\left\{2,\dots,8\right\}$, we independently simulate $\mathbf{X}_{n}$ and $\mathbf{Y}_{n}$ with each $X_{i}$ and $Y_{i}$ ($i\in\left[n\right]$), i.i.d., from the distributions characterized by $f_{1}$ and $h_{1}$, respectively. In E2, we target $f_{2}$ with $h$-MLLEs over the same ranges of $k$ and $n$, but with $h_{2}=\beta\left(\cdot;1,1\right)$–the density of the uniform distribution. For each $k$ and $n$, we simulate $\mathbf{X}_{n}$ and $\mathbf{Y}_{n}$ respectively from distributions characterized by $f_{2}$ and $h_{2}$.
@@
To analyze the sample of $7\times 6\times 50=2100$ observations of relationship between the values $\left(K_{k,n,l}\right)_{l\in\left[r\right]}$ and the corresponding values of $k$ and $n$, we use non-linear least squares (Amemiya, 1985, Sec. 4.3) to fit the regression relationship:
@@
We obtain $95\%$ asymptotic confidence intervals for the estimates of the regression parameters $a_{0},a_{1},a_{2},b_{1},b_{2}\in\mathbb{R}$, under the assumption of potential mis-specification of (16), by using the sandwich estimator for the asymptotic covariance matrix (cf. White 1982). We include the code for these experiments in Appendix B.
@@ 6.3 Results
These observations provide empirical evidence towards the fact that the rate of decrease of $\operatorname*{{\mathrm{\mathbf{E}}}}\left[K_{k,n,l}\right]$ is at least $1/k$ and $1/\sqrt{n}$, respectively, for $k$ and $n$, at least over the simulation scenarios. These fast rates of fit over small values of $n$ and $k$ may be indicative of a diminishing returns of fit phenomenon, as discussed in Cadez and Smyth (2000) or the so-called elbow phenomenon (see, e.g., Ritter 2014, Sec. 4.2), whereupon the rate of decrease in average loss for small values of $k$ is fast and becomes slower as $k$ increases, converging to some asymptotic rate. This is also the reason why we do not include the outcomes when $k=1$, as the drop in $\operatorname*{{\mathrm{\mathbf{E}}}}\left[K_{k,n,l}\right]$ between $k=1$ and $k=2$ is so dramatic that it makes our simulated data ill-fitted by any model of form (16). As such, we do not view Theorem 9 as being pessimistic in light of these phenomena, as it applies uniformly over all values of $k$ and $n$.
@@ 7 Conclusion
The estimation of probability densities using finite mixtures from some base class $\mathcal{P}$ appears often in machine learning and statistical inference as a natural method for modeling underlying data generating processes. In this work, we sought to provide novel generalization bounds for such mixture estimators. To this end, we introduce the family of $h$-lifted KL divergences for densities on compact supports, within the family of Bregman divergences, which correspond to risk functions that can be bounded, even when densities in the class $\mathcal{P}$ are not bounded away from zero, unlike the standard KL divergence. Unlike the least-squares loss, the corresponding maximum $h$-likelihood estimation problem can be computed via an MM algorithm, mirroring the availability of EM algorithms for the maximum likelihood problem corresponding to the KL divergence. Along with our derivations of generalization bounds that achieve the same rates as the best-known bounds for the KL divergence and least square loss, we also provide numerical evidence towards the correctness of these bounds in the case when $\mathcal{P}$ corresponds to beta distribution densities.

### v2

@@ 4 Numerical experiments
Here we discuss the computability and computation of $\mathrm{KL}_{h}$ estimation problems and provide empirical evidence towards the rates obtained in Theorem 5. Specifically, we seek to develop a methodology for computing $h$-MLLEs, and to use numerical experiments to demonstrate that the sequence of expected $h$-lifted KL divergences between some density $f$ and a sequence of $k$-component mixture densities from a suitable class $\mathcal{P}$, estimated using $n$ observations does indeed decrease at rates proportional to $1/k$ and $1/\sqrt{n}$, as $k$ and $n$ increase.
The code for all simulations and analyses in Experiments 1 and 2 is available in both the R and Python programming languages. The code repository is available here: https://github.com/hiendn/LiftedLikelihood.
@@ 4.1 Minorization–Maximization algorithm
One solution for computing (5) is to employ an MM algorithm. To do so, we first write the objective of (5) as
@@
$\lim_{s\rightarrow\infty}\inf_{\psi_{k}^{*}\in\Psi_{k}^{*}}\left\|\psi_{k}^{\left(s\right)}-\psi_{k}^{*}\right\|_{2}=0,\text{ where }\Psi_{k}^{*}=\left\{\psi_{k}^{*}\in\Psi_{k}:\left.\frac{\partial L_{h,n}}{\partial\psi_{k}}\right|_{\psi_{k}=\psi_{k}^{*}}=0\right\}.$
@@ 4.2 Experimental setup
where $\chi_{\mathcal{A}}$ is the characteristic function that takes value $1$ if $x\in\mathcal{A}$ and 00, otherwise. Note that neither $f_{1}$ nor $f_{2}$ are in $\mathcal{C}$. In particular, $f_{1}\left(x\right)=0$ when $x\in\left(\frac{2}{5},\frac{3}{5}\right)$, and $f_{2}(x)=0$ when $x=1/2$, and hence neither densities are bounded away from 0, on $\mathcal{X}$. Thus, the theory of Rakhlin et al. (2005) cannot be applied to provide bounds for the expected KL divergence between MLEs of beta mixtures and the pair of targets. We visualize $f_{1}$ and $f_{2}$ in Figure 1.
@@
To observe the rate of decrease of the $h$-lifted KL divergence between the targets and respective sequences of $h$-MLLEs, we conduct two experiments E1 and E2. In E1, our target density is set to $f_{1}$ and $h_{1}=\beta\left(\cdot;1/2,1/2\right)$. For each $n\in\left\{2^{10},\dots,2^{15}\right\}$ and $k\in\left\{2,\dots,8\right\}$, we independently simulate $\mathbf{X}_{n}$ and $\mathbf{Y}_{n}$ with each $X_{i}$ and $Y_{i}$ ($i\in\left[n\right]$), i.i.d., from the distributions characterized by $f_{1}$ and $h_{1}$, respectively. In E2, we target $f_{2}$ with $h$-MLLEs over the same ranges of $k$ and $n$, but with $h_{2}=\beta\left(\cdot;1,1\right)$–the density of the uniform distribution. For each $k$ and $n$, we simulate $\mathbf{X}_{n}$ and $\mathbf{Y}_{n}$, respectively, from distributions characterized by $f_{2}$ and $h_{2}$.
@@
To analyze the sample of $7\times 6\times 50=2100$ observations of relationship between the values $\left(K_{k,n,l}\right)_{l\in[r]}$ and the corresponding values of $k$ and $n$, we use non-linear least squares (Amemiya, 1985, Sec. 4.3) to fit the regression relationship
@@
We obtain $95\%$ asymptotic confidence intervals for the estimates of the regression parameters $a_{0}$, $a_{1}$, $a_{2}$, $b_{1}$, $b_{2}\in\mathbb{R}$, under the assumption of potential mis-specification of (10), by using the sandwich estimator for the asymptotic covariance matrix (cf. White 1982).
@@ 4.3 Results
These observations provide empirical evidence towards the fact that the rate of decrease of $\operatorname*{{\mathrm{\mathbf{E}}}}\left[K_{k,n,l}\right]$ is at least $1/k$ and $1/\sqrt{n}$, respectively, for $k$ and $n$, at least over the simulation scenarios. These fast rates of fit over small values of $n$ and $k$ may be indicative of a diminishing returns of fit phenomenon, as discussed in Cadez & Smyth (2000) or the so-called elbow phenomenon (see, e.g., Ritter 2014, Sec. 4.2), whereupon the rate of decrease in average loss for small values of $k$ is fast and becomes slower as $k$ increases, converging to some asymptotic rate. This is also the reason why we do not include the outcomes when $k=1$, as the drop in $\operatorname*{{\mathrm{\mathbf{E}}}}\left[K_{k,n,l}\right]$ between $k=1$ and $k=2$ is so dramatic that it makes our simulated data ill-fitted by any model of form (10). As such, we do not view Theorem 5 as being pessimistic in light of these phenomena, as it applies uniformly over all values of $k$ and $n$.
@@ 5 Conclusion
The estimation of probability densities using finite mixtures from some base class $\mathcal{P}$ appears often in machine learning and statistical inference as a natural method for modelling underlying data generating processes. In this work, we pursue novel generalization bounds for such mixture estimators. To this end, we introduce the family of $h$-lifted KL divergences for densities on compact supports, within the family of Bregman divergences, which correspond to risk functions that can be bounded, even when densities in the class $\mathcal{P}$ are not bounded away from zero, unlike the standard KL divergence.
Unlike the least-squares loss, the corresponding maximum $h$-likelihood estimation problem can be computed via an MM algorithm, mirroring the availability of EM algorithms for the maximum likelihood problem corresponding to the KL divergence. Along with our derivations of generalization bounds that achieve the same rates as the best-known bounds for the KL divergence and least square loss, we also provide numerical evidence towards the correctness of these bounds in the case when $\mathcal{P}$ corresponds to beta densities.